Bank Boston has a branch at Bryant College. The branch is busiest at the beginning of the college year when freshmen and transfer students open accounts. This year, freshmen arrived at the office at a rate of 40 per day (assume 8-hour days). On average, it takes the Bank Boston staff person about ten minutes to process each account application. The bank is considering having one or two tellers. Each teller is paid $12 per hour and the cost of waiting in line is assumed to be $8 per hour. What is the total daily waiting cost for the two-teller model?
Answer #1

Solution:
Proportion of time that staffers are bust

= 20/3 /8 = 20/24 = .833


Total wait cost = [(5 * .035 hours) (8)] * 8
= $11.20